### v1

# A Unified Framework for Input Feature Attribution Analysis

Abstract
Explaining the decision-making process of machine learning models is crucial for ensuring their reliability and fairness. One popular explanation form highlights key input features, such as i) tokens (e.g., Shapley Values and Integrated Gradients), ii) interactions between tokens (e.g., Bivariate Shapley and Attention-based methods), or iii) interactions between spans of the input (e.g., Louvain Span Interactions). However, these explanation types have only been studied in isolation, making it difficult to judge their respective applicability. To bridge this gap, we propose a unified framework that facilitates a direct comparison between highlight and interactive explanations comprised of four diagnostic properties [^1] We make the code for the framework available at https://github.com/copenlu/A-unified-framework-for-highlight-explanations .. Through extensive analysis across these three types of input feature explanations–each utilizing three different explanation techniques–across two datasets and two models, we reveal that each explanation type excels in terms of different diagnostic properties. In our experiments, highlight explanations are the most faithful to a model’s prediction, and interactive explanations provide better utility for learning to simulate a model’s predictions. These insights further highlight the need for future research to develop combined methods that enhance all diagnostic properties.
1 Introduction
Figure 1: An example of the three types of Input Feature Explanations on an instance from the SNLI dataset, with their two most important pieces of explanation (token, token tuple, span tuple, correspondingly).
Input feature explanations reveal how a model makes decisions based on the input for which the prediction is made. Among these, Token Explanations (TokenEx; e.g., Shapley Values Lundberg and Lee (2017) and Integrated Gradients (Sundararajan et al., 2017)) are widely used to assign importance scores to specific segments of the input, such as tokens in natural language processing tasks. However, complex reasoning tasks usually require reasoning across multiple pieces of text, e.g., Fact Checking is performed given a claim and evidence, Natural Language Inference is performed given a premise and a hypothesis, where Token Explanations can be insufficient to present the relations employed between the different parts of the input. To this end, Token Interactive Explanations (TokenIntEx; e.g., Bivariate Shapley Masoomi et al. (2023), and Layer-wise Attention Attribution (Ye et al., 2021)) provide importance scores for interactions between two tokens of the input. Further, to enhance the semantic coherence of these explanations, Span Interactive Explanations (SpanIntEx; e.g., Louvain community detection (Ray Choudhury et al., 2023) model interactions across tuples of input spans. An example showcasing these three types of input feature explanations is given in Figure 1.
It is crucial to develop rigorous and comprehensive evaluation frameworks to ensure the principled selection of the most suitable explanation in a practical application Yu et al. (2024) and systematic progress in the development of different types of explanations Atanasova et al. (2022); Jolly et al. (2022). However, these input feature explanations have typically been studied in isolation, with few exceptions, where explanation methods of the same type are compared Atanasova et al. (2020); DeYoung et al. (2020); Janizek et al. (2021); Ray Choudhury et al. (2023). Moreover, evaluations of interactive explanations have been restricted to one property. To address these gaps, we propose a unified framework that facilitates a direct comparison between highlight and interactive input feature explanations on a suite of four diagnostic properties.
Unified Evaluation Framework. Our unified evaluation framework consists of four essential diagnostic properties – Faithfulness, Agreement with Human Annotation, Simulatability, and Complexity. They are the most widely used for Token explanations and include the only property used for interactive explanations – Faithfulness. Faithfulness (Chen et al., 2020, 2021; Ray Choudhury et al., 2023) measures the extent to which explanations accurately reflect the actual reasons used by the model in its predictions. Agreement with Human Annotation (Atanasova et al., 2020) evaluates whether explanations exhibit an inductive bias akin to human reasoning, potentially enhancing their plausibility to end users. Simulatablity (Pruthi et al., 2022) estimates whether the explanations are useful to an agent for replicating the model’s decisions. Finally, Complexity (Bhatt et al., 2021) evaluates whether the explanations are easy to comprehend by presenting distinctive key features rather than a uniform distribution of importance scores. Crucially, we propose extensions of these properties such that they can be applied to different types of input feature explanations – token (TokenEx), two tokens (TokenIntEx), or more than two tokens (SpanIntEx).
Characterising Input Feature Explanations. We conduct an extensive analysis across two different tasks and language models, and three different explanation techniques for each input feature explanation type. Our findings indicate that TokenEx explanations exhibit greater faithfulness than SpanIntEx and TokenIntEx explanations. Additionally, SpanIntEx and TokenIntEx explanations align more closely with human annotations at the token level compared to TokenEx explanations. Moreover, SpanIntEx explanations demonstrate higher interaction overlap with human annotations than TokenIntEx explanations. Further, SpanIntEx explanations are found to be most beneficial for agent simulatability. Finally, our results suggest that SpanIntEx and TokenEx explanations tend to be easier to comprehend.
Overall, our extensive analysis reveals the varied strengths of each explanation type across different diagnostic properties. These insights underscore the importance of developing integrated methods that combine the different types of explanations to enhance all diagnostic properties in future research.
2 A Unified Evaluation Framework for Highlight Explanations
To systematically compare feature attribution explanations of different types, we present a unified framework comprised of four widely employed diagnostic properties: Faithfulness, Agreement with Human Annotations, Simulatability, and Complexity. This section formally introduces these properties and outlines our proposed extensions to allow for their application across different input feature explanation types.
2.1 Preliminaries
We start with a dataset $D$, and a model $M$ fine-tuned on $D$. An instance $x\in D$ is comprised of two parts, e.g., a claim and an evidence, the first consisting of $m$ tokens, and the second – of $n$ tokens. We apply an explanation attribution method $A_{E}$ of type $E\in\{\texttt{TokenEx},\texttt{TokenIntEx},\texttt{SpanIntEx}\}$ to $M$, and each instance $x\in D$: $A_{E}(M,x)=\{e^{x}_{k},a^{x}_{k}|k\in[0,s-1]\}$, where $e^{x}_{k}$ is a span/a pair of spans from the input and $a^{x}_{k}$ denotes its importance score. $k$ is the importance ranking of the corresponding piece of explanation. $s$ is an upper limit for the number of most important pieces of explanation for an instance, such that: $s\in[1,m+n]$ for TokenEx, $s\in[1,m\cdot n]$ for TokenIntEx, and for SpanIntEx $s$ varies for each instance with $s\in[1,f],f<m!\cdot n!$. Depending on the explanation type $E$, $e^{x}_{k}$ can consist of: one token $x_{i}$ for TokenEx, $i\in[0,m+n-1]$, one token pair $(x_{p},x_{q})$ for TokenIntEx, where $p\in[0,m-1]$ and $q\in[m,m+n-1]$, one span pair $((x_{s},\dots,x_{s+l_{1}}),(x_{t},\dots,x_{t+l_{2}}))$ for SpanIntEx, where $s,s+l_{1}\in[0,m-1]$ and $t,t+l_{2}\in[m,m+n-1]$.
Considering a particular threshold for the number of most important explanation pieces $s$, we compute the total set of input tokens involved in the presented explanation for $x$:
$T_{E,M}(x,s)=\{x_{i}|x_{i}\in{e}_{k}^{x},k\in[0,s-1]\}$ (1)
When computing the different diagnostic properties, previous works have adopted a range of thresholds $s$. Moreover, as noted above, the maximum possible value for $s$ can vary across different types of input feature explanations. Additionally, the number of tokens included in the top-$k$ important explanations can differ substantially among explanation types: in top-1 we can have: in a top-1 setting, the explanations consist of 1 token for TokenIntEx, 2 tokens for TokenIntEx, and more than 2 tokens for SpanIntEx Ray Choudhury et al. (2023); DeYoung et al. (2020). This variability makes it difficult to compare results across different explanation types. To address this problem, we propose extensions for each of the studied diagnostic properties that result in unified diagnostic properties which can be used for a direct comparison of the different types of input feature explanations.
2.2 Faithfulness
Faithfulness DeYoung et al. (2020) assesses whether explanations accurately reflect the model’s decision-making process. It involves two aspects – Comprehensiveness and Sufficiency – measured as the number of the model’s prediction changes after keeping (SP) or omitting (CP) $k$ of the most important portions of the input, respectively:
$CP(x,E,k)=\left\{\begin{array}[]{ll}0,&\text{if }f(x-T_{E,M}(x,k))=f(x)\\ 1,&\text{otherwise }\end{array}\right\}$ (2)
$SP(X,E,k)=\left\{\begin{array}[]{ll}1,&\text{if }f(T_{E,M}(x,k))=f(x)\\ 0,&\text{otherwise }\end{array}\right\}$ (3)
To take different thresholds $k$, we average over $k\in[0,s=m+n-1]$. We then also average the results across instances within $D$ to compute the final Comprehensiveness/Sufficiency score:
$Comp_{E}/Suff_{E}=\frac{\sum_{x\in{D}}^{|D|}\sum_{k=0}^{s}CP/SP(x,E,k)}{|D|}$ (4)
Unified Faithfulness. To facilitate the comparison of faithfulness scores, we introduce a dynamic threshold $\theta_{x,k}$. This threshold represents the number of unique tokens used for a perturbation on input $x$, same across all explanation types. Since top-$k$ explanations from SpanIntEx typically contain more tokens than TokenEx or TokenIntEx, the number of tokens in top-$k$ SpanIntEx explanations sets the number of unique tokens used for perturbations across all types of input feature explanations:
$\theta_{x,k}=\left|T_{\texttt{SpanIntEx},M}(x,k)\right|$ (5)
Thus, $\theta_{x,k}$ becomes a dynamic threshold that adapts based on each instance’s top-$k$ SpanIntEx explanation tokens. We then adjust the number of top-$k$ explanations selected from TokenEx and TokenIntEx, $k_{\texttt{TokenEx}}$ and $k_{\texttt{TokenIntEx}}$, correspondingly, to result in the same number of perturbed tokens $\theta_{x,k}$:
$\left|T_{\texttt{TokenEx},M}(x,k_{\texttt{TokenEx}}(x))\right|=\theta_{x,k}$ (6)
$\left|T_{\texttt{TokenIntEx},M}(x,k_{\texttt{TokenIntEx}}(x))\right|=\theta_{x,k}$ (7)
Furthermore, a Random baseline is established, where the tokens for perturbation are selected randomly so that it matches the average $\theta_{x,k}$ across $D$.
2.3 Agreement with Human Annotation
Agreement with Human Annotation has been used to assess the overlap between generated and human-annotated TokenEx explanations Atanasova et al. (2020), which can indicate the plausibility of the generated explanations to end users. For $E=\texttt{TokenEx}$, the measure is computed by calculating a precision score for the top-$k$ most important explanations compared to the gold Human Annotations.
For $E=\texttt{TokenEx}$, $s$ are explanations extracted from instance $x$, where $s=m+n$, $s$ thresholds are set corresponding to each explanation’s attribution score, forming the threshold list $\omega_{E}(x)$:
$\omega_{E}(x)=[a^{x}_{0},...,a^{x}_{s}]$ (8)
By selecting explanations with higher attribution scores than each threshold in $\omega_{E}(x)$, $s$ targeted explanation sets are obtained, where $C_{E}(x,i)$ represents the set for the $i^{th}$ threshold:
$C_{E}(x,i)=\{e^{x}_{j}:a^{x}_{j}<=a^{x}_{i}\}$ (9)
Comparing these sets with the golden explanation set $e^{G}$, $s$ precision-recall pairs can be derived. Average Precision (AP) is then obtained by weighting the precision value with the corresponding recall increase:
$P_{i}/R_{i}(x,e^{G},E)=Pre/Rec(C_{E}(x,i),e^{G})$ (10)
$AP_{E}(x,e^{G})=\sum_{i=0}^{s}(R_{i}-R_{i-1})*P_{i}$ (11)
Finally, Mean AP (MAP) is calculated by averaging AP across the instances in the dataset $D$:
$MAP_{E}=\frac{\sum_{x\in D}^{|D|}AP_{E}(x,e^{G})}{|D|}$ (12)
Unified Agreement with Human Annotation Measure. For a fair comparison between the different types of explanations, the thresholds $\omega_{E}(x)$ for including the same number of tokens across the different types of explanations follows the procedure set in the Unified Faithfulness property (§2.2). In particular, we start with the thresholds set for SpanIntEx, and adjust the thresholds for TokenEx/TokenIntEx to reach the same number corresponding number of tokens. Furthermore, we measure agreement with human annotations at the interaction level for the gold SpanIntEx/TokenIntEx explanations and at the token level for gold TokenEx explanations.
Interaction-level Agreement with Human Annotation.
For a fair comparison between TokenIntEx and SpanIntEx explanations, we adapt $MAP_{\texttt{TokenEx}}$ to the interaction level.
We compute the mean average precision (Eq. 12) w.r.t. the human-annotated TokenIntEx/SpanIntEx sets.
Token-level Agreement with Human Annotation.
For a fair comparison between TokenEx and TokenIntEx/SpanIntEx explanations, we extract tokens from TokenIntEx/SpanIntEx explanations and compare them with tokens from golden TokenIntEx/SpanIntEx explanations. The threshold lists $\omega_{\texttt{TokenIntEx}/\texttt{SpanIntEx}}(x)$ are used, but the targeted sets $C_{E_{token}}(x,i)$ contain tokens extracted from TokenIntEx/SpanIntEx, correspondingly.
The golden set $S_{E_{token}}(x)$ aggregates tokens from golden TokenIntEx/SpanIntEx explanations. The token-level agreement scores $MAP_{\texttt{TokenIntEx}_{token}}/MAP_{\texttt{SpanIntEx}_{token}}$ accordingly, following the procedure set for $MAP_{\texttt{TokenEx}}$ (Eq. 12).
We also set a Random baseline, where the number of randomly selected span pairs, token pairs, or tokens for each instance matches the average number of SpanIntEx explanations per instance across all attribution methods.
2.4 Simulatability
Simulatability was initially proposed to measure how accurately humans can predict a model’s outputs based on its explanations Chen et al. (2023). Previous studies have demonstrated that Simulatability can be approximated using an automated agent model as a surrogate for human understanding Pruthi et al. (2022). Given the established positive correlation between Simulatability and human evaluation of explanation utility from previous research, we integrate the Simulatability scores obtained from the agent model with different explanation types to approximate their utility for humans.
In prior research, an agent model $AM$, sharing the same architecture as the original model $M$, was introduced. $AM$ was trained to simulate $M$’s predictions $Y^{\prime}$ using post-hoc TokenEx explanations. During $AM$’s training phase, the top-$k$ explanations extracted for $M$ were provided as additional signals. In comparison, another agent model, $AM_{O}$, was trained without explanation guidance as a baseline on the same training set. During the testing phase, the simulation accuracy of both $AM$ and $AM_{O}$ over the shared dataset $D$ is calculated. The difference between their accuracies is interpreted as the explanation’s effect in enhancing the original model’s simulatability:
$Sim=ACC(AM(D),Y^{\prime})-ACC(AM_{O}(D),Y^{\prime})$ (13)
Unified Simulatability.
To compare the simulation utility of different explanation types, we train a separate agent model, denoted as $AM_{E}$, for each explanation type $E$, and calculate the corresponding simulation performance on the common test set. For a fair comparison across the different explanation types, we first ensure top-$k_{E}$ explanations are presented for assisting the agent’s training for explanations type $E$, following the same method stated in Section §2.2. This approach ensures that each model is exposed to the same quantity of explanation information, specifically the number of unique tokens from the different explanation types.
During the training phase of the agent model, we introduce the explanations of type $E$ respectively into the learning of $AM_{E}$; we supplement the input sequence $x$ with top-$k_{E}$ explanations instead so that the agent model will be trained with the same mechanism whether the explanations are provided or not, and each training instance will contain the input sequence $x_{E}$ and golden label $Y^{\prime}$ which is predicted by the original model $M$. Specifically, we examine two different ways of presenting the explanations as part of the original input sequence, named $I_{Symbol}$ and $I_{Text}$; all aim to ensure the explanations of different types are inserted similarly. See the details of insertion forms in Appendix A.
At test time, the F1 scores of all agent models – $AM_{E}$ and $AM_{O}$, over the shared dataset $D$ are calculated. Note that the explanations of type $E$ are also provided to the input for the unseen instances for agent model $AM_{E}$. The difference between each agent model $AM_{E}$ and $AM_{O}$ on simulation performance is the final metric indicating how much this specific type of input feature explanation enhances the model’s simulatability, noted as $RSF_{E}$.
$SF_{E}=F1(AM_{E}(x_{E}),Y^{\prime}),x\in D$ (14)
$SF_{O}=F1(AM_{O}(x),Y^{\prime}),x\in D$ (15)
$RSF_{E}=SF_{E}-SF_{O}$ (16)
2.5 Complexity
Feature attribution explanations are designed to aid human understanding of a model’s reasoning over specific instances. However, since humans have a limited capacity to process large amounts of information simultaneously, these explanations need to be easy to comprehend. Even if we select only the top-$k$ attributions with the highest importance scores, they need to be distinctive as opposed to the attribution scores having a uniform distribution.
To measure this complexity of attribution scores, Bhatt et al. (2021) employ entropy Rényi (1961) over the attribution scores of all the produced highlight explanations. Higher entropy means different features have similar attribution scores, where the simplest explanation, with low entropy, would be concentrated on one feature. The calculation process is noted as follows, where m+n is the total number of generated TokenEx explanations, and all explanations are considered for the complexity score computation.
$P(x,p)=\frac{\left|a^{x}_{p}\right|}{\sum_{q=1}^{m+n}\left|a^{x}_{q}\right|}$ (17)
$CL(X)=-\sum_{p=1}^{m+n}P(x,p)\ln(P(x,p))$ (18)
Unified Complexity. Building on their approach, we also employ Entropy to measure the complexity of explanation attribution scores across different types of input feature explanations. To ensure a fair comparison, we maintain consistency in the size of the chosen explanation lists across all types for the same instance, denoted as $k_{x}$, as the number of generated TokenEx/TokenIntEx/SpanIntEx explanations vary for the same instance. The complexity score of the top-$k_{x}$ explanation list under type $E$, denoted as $CL_{E}(x,k_{x})$, is calculated as:
$P_{E}(x,k_{x},i)=\frac{\left|a^{x}_{i}\right|}{\sum_{j=1}^{k_{x}}\left|a^{x}_{j}\right|}$ (19)
$CL_{E}(x,k_{x})=-\sum_{i=1}^{k_{x}}P_{E}(x,k_{x},i)\ln(P_{E}(x,k_{x},i))$ (20)
where $a^{x}_{i}$/$a^{x}_{j}$ represent the $i$/$j$th highest attribution score from the explanation set for $x$. $P_{E}(x,k_{x},i)$ is the fraction of the attribution score of the $i^{th}$ most important explanation compared to the sum of the scores within all the top $k_{x}$ chosen explanations.
Finally, by applying this calculation of the complexity score regarding explanation attribution lists for type $E$ to all instances, we obtain the complexity score of each explanation type as $CL_{E}$.
$CL_{E}=\frac{\sum_{x\in D}^{\left|D\right|}CL_{E}(x,k_{x})}{\left|D\right|}$ (21)
Notably, we calculate $k_{x}$ using the number of SpanIntEx explanations within the instance, as the number of explanations of this type is unknown until they are generated, as discussed in §2.1.
3 Experimental Setup
3.1 Datasets
We select the natural language inference dataset SNLI (Bowman et al., 2015), where instances consist of a premise, a hypothesis, and a label $y\in\{entailment,neutral,contradiction\}$. Additionally, we select the fact checking dataset FEVER (Thorne et al., 2018), where instances consist of a claim, an evidence, and a label $y\in\{entailment,neutral,contradiction\}$. For generating the input feature explanations, we sample 4000 instances from each train, dev, and test set due to the computational requirements especially pronounced for Shapley-based explanations Atanasova et al. (2020). To evaluate agreement with human annotations, we describe in Appendix B the datasets used, which contain various types of input feature explanations for SNLI and FEVER annotated by humans.
3.2 Input Feature Explanation Methods
To generate importance scores, as techniques for TokenEx we select three common explanation techniques – Shapley (Lundberg and Lee, 2017), Attention (DeYoung et al., 2020), and Integrated Gradients (IG, Sundararajan et al. (2017)). For TokenIntEx, we employ Bivariate Shapley Masoomi et al. (2023), Attention (Clark et al., 2019), and Layer-wise Attention Attribution (Ye et al., 2021). Following Ray Choudhury et al. (2023), we apply the Louvain algorithm (Blondel et al., 2008) on top of each of the three selected TokenIntEx to generate importance scores for SpanIntEx, where the importance score of each span interaction is averaged over the importance scores of the token interactions within it.
Notably, the techniques applied for generating TokenIntEx are the bivariate version of the techniques used for generating TokenEx; e.g. Layer-wise Attention Attribution uses IG. Furthermore, SpanIntEx are generated from importance scores of TokenIntEx. We will refer to Shapley, Attention, and IG as the explanation base types used for generating all types of input feature explanations for brevity. See the details of generating all input feature explanations with different explainability techniques in Appendix D.
3.3 Models
We use the BERT-base-uncased model (Devlin et al., 2019) with 12 encoder layers, and the BART-base-uncased model (Lewis et al., 2020) with 6 encoder and 6 decoder layers, as common representatives of the encoder and the encoder-decoder Transformer architecture Vaswani et al. (2017). Our model choice is particularly influenced by the substantial computational requirements of the input feature explanations, especially pronounced for Shapley Atanasova et al. (2020). Additionally, our choice is guided by the need to directly access the models’ internals for generating IG and Attention-based explanations. Furthermore, while our framework currently utilizes the said models, it is designed to be easily adaptable to other models or newly developed explainability techniques, provided that there are more robust computational resources available.
We fine-tune the base models on SNLI and FEVER and use them to generate explanations. Their performance on the corresponding test sets is shown in Appendix C. For the Simulation property, we follow existing work Fernandes et al. (2022); Pruthi et al. (2022) and train simulator agent models (see §2.4) with the same architectures as the base ones. Following Fernandes et al. (2022), we split the test set into train/dev/test for the training of the agent model that simulates the original model’s predictions.
4 Results and Discussion
(a) SNLI dataset, BERT model.
(b) SNLI dataset, BART model.
(c) FEVER dataset, BERT model.
(d) FEVER dataset, BART model.
Figure 2: Unified evaluation framework (§2) results for all feature attribution methods (§3.2).
We now present the results of our unified evaluation framework (§2) illustrated in Figure 2. Our results include span the input feature explanations SpanIntEx, TokenIntEx, and TokenEx  (§3.2), two models (§3.3), two datasets (§3.1), and three base explanation techniques per input feature explanation type (§3.2). For Simulatability, we select the results of $I_{sym}$, as this form avoids repeating the input text and increasing the input size substantially. For Agreement with Human Annotation, we select the results of Token-level agreement as they are present for all types of explanations. Appendix E lists all detailed results per property.
4.1 Faithfulness
Unified Comprehensiveness. Across both datasets and models, TokenEx and TokenIntEx are identified as the most comprehensive explanation types, achieving the highest scores in 7/12 and 5/12 cases, respectively. SpanIntEx, designed to enhance the semantic coherence of interactive explanations by including additional context, often incorporates tokens that do not directly contribute to the model’s prediction, thus explaining its lower comprehensiveness scores. Compared to the random baseline, TokenEx and SpanIntEx always outperform it, while TokenIntEx mostly underperforms it when based on IG. Across the base explanation techniques, TokenEx performs best when based on Attention for BERT and on Shapley for BART, indicating that different base explanation techniques can perform better for different architectures. Both TokenIntEx and SpanIntEx show optimal performance when based on Shapley and Attention. Overall, the results indicate a stronger performance of Attention and Shapley over IG across all explanation types.
Unified Sufficiency. SpanIntEx ranks as the most sufficient explanation type in 7/12 cases, surpassing TokenEx, which performs well in only 3/12 cases. While contrary to SpanIntEx Comprehensiveness performance, we attribute this to the semantic coherence of the extracted top spans, which provide more meaningful information. Unlike TokenEx, which consistently outperforms the random baseline, TokenIntEx and SpanIntEx struggle to outperform it on FEVER, likely due to the longer input length compared to SNLI, posing more challenges for the explanations to accurately unveil the model’s internal processes. The results from different base explanation techniques show no clear trends, indicating a significant variability stemming from the specific dataset and model architecture.
4.2 Agreement with Human Annotation
In comparing token-level agreement, SpanIntEx and TokenIntEx show higher agreement scores with human annotations than TokenEx. Similarly, we find that SpanIntEx consistently achieves higher agreement with human interaction-level annotators, especially when based on Attention scores (see Appendix E). This indicates that SpanIntEx is more plausible to humans than the other types of explanations due to their enhanced semantics coherence. In contrast, TokenEx often scores lower than the random baseline. Moreover, considering SpanIntEx’s lower performance in terms of Comprehensiveness, there emerges a distinct trade-off between comprehensiveness and plausibility. Across the base explanation techniques, IG performs best for the FEVER dataset; IG and Attention perform best for SNLI. In addition, we find that for the interaction-level agreement, TokenIntEx and SpanIntEx perform worst when based on Shapley. The lower agreement results for Shapley compared to its better comparative performance on Comprehensiveness, once again indicate an existing trade-off between the two properties.
4.3 Simulatability
Our results show that SpanIntEx achieves the highest simulatability in 9/12 cases, helping the agent model accurately reproduce the original model’s predictions. This again underscores the critical role of contextual information and enhanced semantic coherence provided by SpanIntEx explanations. Notably, providing SpanIntEx explanations to agents improves their ability to simulate the original model by up to 7.9 F1 points compared to scenarios without explanations. Among the base explanation techniques, IG consistently performs best for SpanIntEx; other techniques do not exhibit a clear trend.
4.4 Complexity
SpanIntEx and TokenEx generally achieve similar complexity scores, which consistently remain lower than those of TokenIntEx. This suggests that TokenEx and SpanIntEx generate more distinctive attribution scores, potentially making them easier for humans to understand. Regarding the base explanation techniques, Attention consistently yields the best complexity scores for BERT across all explanation types. There is however no clear trend for BART. Additionally, as detailed in Appendix E, TokenIntEx frequently underperforms compared to a random baseline, highlighting its complexity.
4.5 Overall
In sum, we find that while TokenEx and TokenIntEx generally provide more comprehensive insights, SpanIntEx performs better in Sufficiency due to its enhanced semantic coherence, and despite occasionally including irrelevant tokens (§4.1). We find that there exists a trade-off between comprehensiveness and the plausibility of explanations (§4.2), suggesting that the most faithful explanations, and in turn, the model’s actual inner workings can be less plausible to end users.
Furthermore, SpanIntEx significantly helps for Simulatability, allowing agents to accurately replicate model decisions (§4.3), which is crucial for practical applications. The Complexity analysis (§4.4) shows that SpanIntEx and TokenEx are easier to comprehend compared to TokenIntEx, underscoring the importance of selecting explanation techniques that cater to the specific needs of the dataset and model architecture.
5 Related Work
Input Feature Explanations. Considerable research exists on extracting explanations from input data. Methods like perturbation-based attribution (e.g., Shapley (Lundberg and Lee, 2017)), attention-based methods (e.g., Attention (Jain and Wallace, 2019; Serrano and Smith, 2019)), and gradient-based methods (e.g., Integrated Gradients (Sundararajan et al., 2017; Serrano and Smith, 2019)) are prevalent for highlighting individual tokens. As individual tokens lack interaction information and thus might be insufficient to explain the model, many attribution methods have been extended to bivariate forms (Masoomi et al., 2023; Janizek et al., 2021; Sundararajan et al., 2017; Ye et al., 2021) to capture input token interactions. More recent work has explored how groups of tokens collectively contribute to model reasoning (Ray Choudhury et al., 2023; Chen et al., 2021). Unlike other work where token groups might consist of tokens from arbitrary positions, Ray Choudhury et al. (2023) introduces a method to explicitly capture span interactions, enhancing the comprehensiveness of explanations by containing the entire spans.
Explanation Evaluation. For evaluating TokenEx explanations, DeYoung et al. (2020); Atanasova et al. (2020) propose metrics to measure how faithful explanations are to the model’s inner reasoning. They also propose to assess the plausibility of explanations to humans by measuring the agreement of TokenIntEx with human annotations. To assess the utility of explanations to humans, Pruthi et al. (2022) propose to use an agent model as a proxy for humans and evaluate whether explanations aid in model simulatability. Complexity Bhatt et al. (2021) measures the distribution of attribution scores of TokenEx and assesses whether the key tokens in token explanations are easily comprehensible to humans. To evaluate TokenIntEx  most works adopt the faithfulness or axiomatic/theoretical path  Tsang et al. (2020); Sundararajan et al. (2020); Janizek et al. (2021). Current work on evaluating SpanIntEx explanations has primarily focused on faithfulness Ray Choudhury et al. (2023). However, since SpanIntEx, TokenIntEx, and TokenEx explanations contain varying amounts of tokens, which e.g. affects the faithfulness test, this makes direct comparisons between different explanation types using existing metrics challenging. To our knowledge, no prior paper has involved all types of input feature explanations within a unified evaluation framework.
6 Conclusion
We introduced a unified evaluation framework for input feature attribution analysis to guide the principled selection of the most suitable explainability technique in practical applications. Our analysis outlines the diverse strengths and trade-offs among TokenEx, TokenIntEx, and SpanIntEx. Our findings particularly underscore SpanIntEx’s superior agreement with human inductive biases and its enhancement of Simulatability, despite its lower comprehensiveness in capturing the model’s reasoning compared to TokenEx and TokenIntEx. This study emphasizes the importance of selecting explanation techniques tailored to specific datasets and model architectures. Future efforts should focus on developing combined methods that enhance both the faithfulness and user comprehensibility of explanations.
Limitations
Our work introduces a unified framework to evaluate input feature explanations across four key properties. We generated three types of explanations using three attribution methods on two transformer models (BERT-base and BART-base) for two NLU tasks (NLI and fact-checking). This enabled us to assess and compare the properties of each explanation type. Due to computational resource limitations, we did not include larger decoder-only models in our evaluation. Future research could explore these models to provide additional insights. Additionally, examining the properties of different explanations in various tasks beyond NLI and fact-checking would be valuable.
Our findings indicate that span interactive explanations (SpanIntEx) have a notable advantage over other types in Agreement with Human Annotation, Simulatability, and Complexity, suggesting they are easier for humans to understand. This insight could inspire future work to leverage SpanIntEx as the input feature explanation in HCI models. However, SpanIntEx shows low comprehensiveness in faithfulness evaluations. The Louvain algorithm, used for SpanIntEx generation, may limit its comprehensiveness despite using different attribution methods for TokenIntEx. Future work should explore better methods for capturing span interactions and possibly combine SpanIntEx and TokenEx for higher faithfulness, as TokenEx demonstrates a stable advantage in comprehensiveness.
Another core finding is that the attribution method significantly affects most diagnostic properties of all explanation types, such as sufficiency. No single attribution method consistently excels across all properties, highlighting the need for continuous evaluation and improvement in attribution methods, particularly for SpanIntEx.
Acknowledgements
[Uncaptioned image][Uncaptioned $\end{array}$ This research was co-funded by the European Union (ERC, ExplainYourself, 101077481), by the Pioneer Centre for AI, DNRF grant number P1, as well as by The Villum Synergy Programme (grant number 40543). Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council. Neither the European Union nor the granting authority can be held responsible for them.
Appendix A Detailed Explanation Insertion Method
To enable a fair comparison among different input feature explanations in terms of simulatability (§2.4), we applied consistent insertion formats to combine the explanations with the original input for training the agent models. This design aims to minimize noise from insertion format differences. We tested two ways, each applicable to all types of input feature explanations, to construct input sequences with inserted explanations of type $E$. These input sequences are denoted $x_{E}$ in §2.4, omitting specific insertion format details for brevity.
For Symbol-Insertion $I_{Symbol}$, we preserve the original input sequence but insert special symbols $<$ and $>$ to quote the tokens (for TokenEx and TokenIntEx) or spans (for SpanIntEx) within the input. Additionally, for TokenEx, we append a ranking mark after each quoted token based on their attribution scores, ranked in descending order. For TokenEx and SpanIntEx, each quoted token/span is also assigned a ranking mark indicating the rank of their respective interactions by attribution score, ensuring tokens/spans from the same interaction share the same mark. This method allows us to generate input sequences combined with different input feature explanations in a consistent symbol insertion format.
For Text-Insertion $I_{Text}$, we append tokens, token tuples, or span tuples to the end of the original input sequence for each explanation type. They are added in the order ranked by descending attribution score. Specifically, for TokenEx, tokens from different TokenExexplanations are separated by semicolons. For TokenIntEx and SpanIntEx, tokens/spans within each interaction are connected by a comma, and different interactions are separated by semicolons. This approach constructs input sequences combined with each type of input feature explanation in a consistent text insertion format.
Appendix B Agreement Dataset Details
To assess how different types of input feature explanations overlap with human annotations, we collected golden explanations of various types from e-SNLI and SpanEx for instances within the SNLI and FEVER datasets, respectively. Detailed information about the annotated explanation types and the number of instances with labeled explanations for these datasets is shown in Table 1. For the SNLI dataset, e-SNLI provides TokenEx explanations, while SpanEx-SNLI includes SpanIntEx explanations and token-level interactions (TokenIntEx explanations). We selected 3,865 overlapping instances and evaluated the human agreement score for different types of input feature explanations. For the FEVER dataset, SpanEx-FEVER includes SpanIntEx and token-level interactions (TokenIntEx explanations). Since no TokenEx explanations are provided, we extracted tokens from the golden TokenIntEx explanations in SpanEx-FEVER as an approximation. These selected instances are also used when evaluating other properties of input feature explanations.
$D$ $E$ Size
SNLI -
549367 Train
9842 Dev
9824 Test
e-SNLI TokenEx
549367 Train
9842 Dev
9824 Test
SpanEx-SNLI
SpanIntEx
TokenIntEx
3865 Test
FEVER -
145449 Train
$9$ Dev
$9$ Test
SpanEx-FEVER
SpanIntEx
TokenIntEx
3206 Test
Table 1: Overview of datasets SNLI (Bowman et al., 2015), FEVER (Thorne et al., 2018), SpanEx (Ray Choudhury et al., 2023) and e-SNLI (Camburu et al., 2018). SpanEx contains instances from SNLI and FEVER datasets, annotated with SpanIntEx explanations including token-level explanations (TokenIntEx explanations). e-SNLI contains instances from SNLI dataset, annotated with TokenEx explanations.
Appendix C Base Model Performance.
As shown in Table 2, we report the performance of fine-tuned BERT-base and BART-base models on the SNLI and FEVER datasets, respectively. These models, fine-tuned for their specific tasks, are used to generate various input feature explanations through different explainability techniques. Importantly, these are the original models that the agent models, as described in §2.4, learn to simulate.
$Model$ F1 score
Dev Test
BERT-SNLI 87.21 88.43
BART-SNLI 86.81 85.40
BERT-FEVER 86.21 89.49
BART-FEVER 85.19 84.88
Table 2: The performance of our BERT-base and BART-base models fine-tuned on SNLI and FEVER datasets, respectively, regarding F1 score(%).
Appendix D Explainability Techniques
In this section, we detail the explainability techniques employed to generate various types of input feature explanations. As outlined in §3.2, we categorize these techniques based on the method used for generating TokenEx, while TokenIntEx explanations stem from their bivariate variants, forming the basis for SpanIntEx explanations.
Shapley. To assign importance scores to each token within the input, we employ the SHAP method following Lundberg and Lee (2017), utilizing Kernel SHAP to approximate Shapley values. For token interactions, we apply Bivariate Shapley  (Masoomi et al., 2023) to assess a token’s influence to model prediction in the presence of another token, particularly when they originate from different parts of the input. The importance score for each token interaction is derived by averaging these mutual influence scores between the involved tokens. With these importance scores for all possible token interactions, we apply the Louvain algorithm  (Ray Choudhury et al., 2023) to identify clusters of neighboring token interactions characterized by dense intra-cluster and sparse inter-cluster relationships, which collectively form span interactions. The importance score for each span interaction is calculated by averaging the importance scores of the token interactions it comprises.
Attention. For each token within the input sequence, we use the self-attention weights between this token and the first token as an indicator of its importance score (Jain and Wallace, 2019). We follow Ray Choudhury et al. (2023) to select the most important attention head in the last layer of the model to obtain these attention weights. For each possible token interaction, we use the method by Clark et al. (2019) to extract and average the attention weights between token pairs from different parts of the input to derive their importance scores, also from the most important head of the last layer. To obtain span interactions and assign them importance scores, we apply the same method to these token interaction scores as described above.
Integrated Gradients. To calculate the importance score for each token in the input sequence, we integrate the gradients of the model’s output with respect to each token embedding, following Sundararajan et al. (2017). For generating the importance scores of token interactions, we use Layer-wise Attention Attribution (Ye et al., 2021), which attributes attention links between pairs of tokens within attention maps with a mechanism similar to Integrated Gradients. These attribution maps are created for each model layer and then aggregated across layers to form a final attribution map. The importance score for each token interaction is calculated as the average value from this final attribution map between the involved tokens. For span interactions, we generate and assign importance scores using the same approach based on the importance scores of the token interactions.
Appendix E Detailed Experiment Results
E.1 Faithfulness
(a) Comprehensiveness on SNLI dataset, the higher the better.
(b) Comprehensiveness on FEVER dataset, the higher the better.
(c) Sufficiency on SNLI dataset, the lower the better.
(d) Sufficiency on FEVER dataset, the lower the better.
Figure 3: Unified Comprehensiveness and Sufficiency of three types of feature attribution explanations on SNLI and FEVER datasets using the BERT model. Subfigures (a) and (c) show Unified Comprehensiveness results, while (b) and (d) show Unified Sufficiency results. Explanations are generated by Shapley (Shap), Attention (Att), and Integrated Gradients (IG). Randomly selected span pairs, token pairs, and tokens are baselines corresponding to explanation type SpanIntEx, TokenIntEx, and TokenEx  and form the group Random baseline (Rand). We set $k=3$ for top span interactions and adjust token counts as per §2.2, also ensuring the random baseline matches the average token count of the top $k$ span interactions.
(a) Comprehensiveness on SNLI dataset, the higher the better.
(b) Comprehensiveness on FEVER dataset, the higher the better.
(c) Sufficiency on SNLI dataset, the lower the better.
(d) Sufficiency on FEVER dataset, the lower the better.
Figure 4: Unified Comprehensiveness and Sufficiency of three types of feature attribution explanations on SNLI and FEVER datasets using the BART model. Subfigures (a) and (c) show Unified Comprehensiveness results, while (b) and (d) show Unified Sufficiency results. Explanations are generated by Shapley (Shap), Attention (Att), and Integrated Gradients (IG). Randomly selected span pairs, token pairs, and tokens are baselines corresponding to explanation type SpanIntExTokenIntEx  and TokenEx  and form the group Random baseline (Rand). We set $k=3$ for top span interactions and adjust token counts as per section §2.2, ensuring the random baseline matches the average token count of the top $k$ span interactions.
E.2 Agreement with Human Annotation
$\mathbf{E}$ Shap Att IG Rand
Interaction level agreement
SpanIntEx 30.18 57.40 39.40 33.82
TokenIntEx 29.02 37.02 35.06 23.42
TokenEx - - - -
Token level agreement
SpanIntEx 75.63 78.26 76.52 76.96
TokenIntEx 74.89 79.60 73.19 74.96
TokenEx 75.54 77.62 70.74 76.33
Table 3: Human Annotation Agreement Results (see §2.3) on SNLI dataset when explanations are generated based on BERT. Interaction-level and Token-level agreement scores, Average Precision(%), compared to human annotations for explanation types SpanIntEx, TokenIntEx, TokenEx generated by Shapley(Shap), Attention(Att), Integradiant Gradients(IG) respectively. Using the same attribution method, the highest alignment score for each category is highlighted in bold. Rand indicates the random baseline as described in §2.3.
$\mathbf{E}$ Shap Att IG Rand
Interaction level agreement
SpanIntEx 19.92 28.12 27.45 19.33
TokenIntEx 3.96 10.27 21.30 10.23
TokenEx - - - -
Token level agreement
SpanIntEx 66.95 68.71 72.24 67.5
TokenIntEx 66.90 67.29 70.50 65.86
TokenEx 58.07 56.88 61.07 63.10
Table 4: Human Annotation Agreement Results (see §2.3) on the FEVER dataset when explanations are generated based on BERT. The rest of the settings are the same as Table 3.
There is a notable gap between interaction-level and token-level agreement scores. For example, in Table 3, the highest interaction-level agreement score for SpanIntEx explanations is 57.40%, while the highest token-level agreement score for SpanIntEx is 78.26%. A similar pattern is observed for TokenIntEx. This suggests that although SpanIntEx and TokenIntEx explanations align more with human reasoning than TokenEx explanations, pairing important tokens or spans into interactions that are plausible to humans remains challenging.
$\mathbf{E}$ Shap Att IG Rand
Interaction level agreement
SpanIntEx 37.36 47.33 34.18 28.25
TokenIntEx 32.36 35.17 33.06 13.80
TokenEx - - - -
Token level agreement
SpanIntEx 80.16 76.28 82.76 70.04
TokenIntEx 84.9 75.92 86.06 75.32
TokenEx 65.74 73.71 70.44 73.34
Table 5: Human Annotation Agreement Results (see §2.3) on SNLI dataset when explanations are generated based on BART. The rest settings are the same as Table 3.
$\mathbf{E}$ Shap Att IG Rand
Interaction level agreement
SpanIntEx 22.86 20.66 18.72 16.76
TokenIntEx 4.71 2.64 10.54 8.64
TokenEx - - - -
Token level agreement
SpanIntEx 68.77 65.33 70.22 69.51
TokenIntEx 67.01 63.40 70.98 68.11
TokenEx 59.43 52.15 57.56 60.93
Table 6: Human Annotation Agreement Results (see §2.3) on FEVER dataset when explanations are generated based on BART. The rest settings are the same as Table 3.
E.3 Simulatability
Regarding insertion formats, for BERT models, text insertion ($I_{Text}$), which adds explanation text to the end of the input sequence, consistently outperforms symbol insertion ($I_{Sym}$), where symbols are added to the original input sequence, as shown in Tables 7 and 8. However, the opposite effect is observed for BART models, as shown in Tables 9 and 10. This indicates that simulatability results are sensitive to the explanation insertion form, highlighting the need for consistency in insertion form when comparing different explanation types.
$\mathbf{D}$ $\mathbf{E}$ Shap Att IG
$\mathbf{SF}$ $\mathbf{RSF}$ $\mathbf{SF}$ $\mathbf{RSF}$ $\mathbf{SF}$ $\mathbf{RSF}$
SNLI SpanIntEx 87.9 3.2 86.7 2.0 88.9 4.2
TokenIntEx 86.6 1.9 85.3 0.6 85.8 1.1
TokenEx 87.4 2.7 85.7 1.0 86.0 1.3
FEVER SpanIntEx 83.9 3.8 85.3 5.2 85.2 5.1
TokenIntEx 82.7 2.6 84.0 3.9 84.4 4.3
TokenEx 84.0 3.9 82.3 2.2 81.8 1.7
Table 7: Simulatability results on SNLI and FEVER with BERT as the model used for all explanations $E\in{\texttt{SpanIntEx},\texttt{TokenIntEx},\texttt{TokenEx}}$ generation with attribution method Shapley, Attention, and Integrated Gradients respectively. Note that insertion form $I_{Sym}$ is adopted for combining the explanations and the original input sequence for agent model $AM_{E}$, as depicted in §2.4. The agent models used for baseline $AM_{O}$, trained without explanations, have simulation F1 scores, as denoted in §2.4, of 84.7% and 80.0% on test set shared with other agent models $AM_{E}$, as denoted in §2.4. We set $k=1$ for top SpanIntEx and calculated the number of top TokenIntEx and TokenEx accordingly as stated in §2.2. The largest increases are highlighted in bold for each dataset with the identical attribution method.
$\mathbf{D}$ $\mathbf{E}$ Shap Att IG
$\mathbf{SF}$ $\mathbf{RSF}$ $\mathbf{SF}$ $\mathbf{RSF}$ $\mathbf{SF}$ $\mathbf{RSF}$
SNLI SpanIntEx 87.8 3.1 87.1 2.4 88.2 3.5
TokenIntEx 86.5 1.8 87.8 3.1 86.4 1.7
TokenEx 87.0 2.3 86.3 1.6 88.4 3.7
FEVER SpanIntEx 85.7 5.6 85.1 5.0 86.0 5.9
TokenIntEx 81.9 1.8 85.6 5.5 84.3 4.2
TokenEx 85.8 5.7 84.5 4.4 82.0 1.9
Table 8: Simulatability results on SNLI and FEVER with BERT as the model used for all input feature explanation generation. Note that insertion form $I_{Text}$ is adopted for combining the explanations and the original input sequence for agent model $AM_{E}$, as depicted in §2.4. The agent models used for baseline $AM_{O}$, which are trained without explanations, have simulation F1 scores of 84.7% and 80.0% on the test sets shared with agent model $AM_{E}$. The other setting is the same as Table 7
$\mathbf{D}$ $\mathbf{E}$ Shap Att IG
$\mathbf{SF}$ $\mathbf{RSF}$ $\mathbf{SF}$ $\mathbf{RSF}$ $\mathbf{SF}$ $\mathbf{RSF}$
SNLI SpanIntEx 87.8 7.9 87.3 7.4 86.5 6.6
TokenIntEx 83.5 3.6 84.2 4.3 84.6 4.7
TokenEx 88.2 8.3 81.4 1.5 85.8 5.9
FEVER SpanIntEx 80.6 7.2 76.1 2.7 75.2 1.8
TokenIntEx 78.9 5.5 75.9 2.5 74.7 1.3
TokenEx 80.1 6.7 75.0 1.6 74.1 0.7
Table 9: Simulatability results on SNLI and FEVER with BART as the model used for all input feature explanation generation. Note that insertion form $I_{Sym}$ is adopted for combining the explanations and the original input sequence for agent model $AM_{E}$, as depicted in §2.4. The base agent models $AM_{O}$ trained without explanations have the simulation f1 scores of 79.9% and 73.4%, respectively on the test sets sharing with other agent models $AM_{E}$. The other setting is the same as Table 7
$\mathbf{D}$ $\mathbf{E}$ Shap Att IG
$\mathbf{SF}$ $\mathbf{RSF}$ $\mathbf{SF}$ $\mathbf{RSF}$ $\mathbf{SF}$ $\mathbf{RSF}$
SNLI SpanIntEx 86.8 6.9 85.0 5.1 84.3 4.4
TokenIntEx 81.2 1.3 82.6 2.7 81.6 1.7
TokenEx 83.3 3.4 83.8 3.9 82.2 2.3
FEVER SpanIntEx 78.2 4.8 75.3 1.9 74.6 1.2
TokenIntEx 74.8 1.4 74.1 0.7 73.9 0.5
TokenEx 77.6 4.2 74.9 1.5 73.6 0.3
Table 10: Simulatability results on SNLI and FEVER with BART as the model used for all input feature explanation generation respectively. Note that insertion form $I_{Text}$ is adopted for combining the explanations and the original input sequence for agent model $AM_{E}$, as depicted in §2.4. The base agent models $AM_{O}$ trained without explanations have the simulation f1 scores of 79.9% and 73.4%, respectively. The other setting is the same as Table 7
E.4 Complexity
Dataset $E$ Shapley Attention IG R U
SNLI SpanIntEx 2.05 1.11 2.10 2.19 2.62
TokenIntEx 1.72 1.43 2.30 - -
TokenEx 2.08 1.64 1.91 - -
FEVER SpanIntEx 2.78 2.22 2.90 2.98 3.18
TokenIntEx 3.12 3.07 3.15 - -
TokenEx 2.76 2.22 2.57 - -
Table 11: Complexity results on SNLI and FEVER datasets for three types of explanations generated by different attribution methods based on BERT model. The Random baseline represents the complexity score obtained by randomly generated scores in the range [0,1], ensuring the same number of scores as the number of explanations used. The Upperbound is calculated by setting all the attribution scores to the same value while ensuring the same number of scores as the number of explanations used. The lowest complexity score for each specific explanation type compared is highlighted in bold when the explanations are generated by each attribution method.
Dataset $E$ Shapley Attention IG R U
SNLI SpanIntEx 1.90 1.93 1.63 2.36 2.76
TokenIntEx 2.50 2.53 2.13 - -
TokenEx 1.86 1.95 1.94 - -
FEVER SpanIntEx 2.73 3.03 2.38 3.13 3.38
TokenIntEx 3.30 3.36 2.93 - -
TokenEx 2.82 3.07 3.19 - -
Table 12: Complexity results on SNLI and FEVER datasets for three types of explanations generated by different attribution methods based on BART model. The other settings are the same as Table 11.